# Preprint submitted to Journal

### v4

@@ Abstract
This paper proposes an algebraic framework for analyzing event execution intervals and sequences, introducing the concept of Optional Intervals Event (OIE) as a core abstraction. $\mathrm{OIE}$ maps real-world events into a 4-tuple structure that stores all feasible execution intervals and sub-event relationships. Based on this, two sequential operations are defined: Complete Sequential Addition ($\oplus$) models concurrent events with a certain degree of equal opportunity within a shared time domain, Complete Sequential Multiplication ($\otimes$) models sequential events with a strict order. And the algebraic properties and orbit spaces of these operations are analyzed.
Addressing the traditional challenge of “simultaneous start”, this paper argues that pointwise temporal equality of timestamps is an unattainable ideal due to physical limits of observational precision. Instead, using sequential operation, “simultaneous start” is reformulated as equal opportunity within an equipotential time window, rather than observed instantaneous equality. This paradigm shift moves the perspective from passive observation to active, constraint-aware planning, providing a more rigorous formal foundation for event scheduling, physical simulation, and computational modeling that can handle inherent observational uncertainties.
@@ 1. How can we accurately describe the so-called “simultaneous start of two events”?
As of the end of 2024, the error of the most advanced atomic clock in the world is $8\times 10^{-19}$ seconds(see Aeppli et al. (2024)), that is to say, it cannot be guaranteed that the so-called simultaneously starting moments of the two observed events can be accurate to $10^{-1000}$ seconds. So, strictly speaking, this scenario is not truly “at the same time”, we should pursue a more rigorous expression. Can we use an operation *, represented by the following expression
@@
In a 100 metres race, there are usually 8 athletes competing, it is a head-to-head type of race. In this paper, we use a finitary operation $\oplus$” to represent a race in which all athletes participate together. Suppose the 8 athletes are arranged according to the lane order, with the $i$-th athlete $athlete_{i}$ in the $i$-th lane. Then a 100 metres race can be expressed as
@@
Swapping the lanes does not affect the starting method of each athlete, they all start running after hearing the starting gun. So intuitively, the operation $\oplus$” satisfies the commutative law.
Downhill skiing is a winter sport. Each skier competes on a track in a certain order one by one, and there can only be one person on the track at a time, so this is a sequential type of race. We use a finitary operator $\otimes$” to represent a race in which all skiers participate. Then a race with 6 skiers can be expressed as
@@
The skier represented on the left side of parameters starts the race earlier than the skier represented on the right side. Intuitively, the operation $\otimes$” does not satisfy the commutative law.
@@
This paper introduces a paradigm shift towards the subjective planning of events. Employing mathematical frameworks from set theory and group theory, it establishes a novel formal methodology for the analysis of event execution and object motion. The subsequent sections are structured as follows:
@@
  • Section 3: Optional Intervals Event (OIE) – Proposes the core abstraction of an Optional Intervals Event, a 4-tuple $(\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A})$ that maps to a real-world event. It stores all feasible execution intervals and the relationships among sub-events, serving as a formal tool for pre-execution planning. The concepts of atomic (AtomOIE) and composite (CompOIE) OIE are defined.
  • Section 4: Two Sequential Operations – Introduces two n-ary finitary operations on sets of OIE instances. Complete Sequential Addition $(\oplus)$ models concurrent events where all participants have a certain degree of equal opportunity to start and end within a specified domain. Complete Sequential Multiplication $(\otimes)$ models strictly ordered sequential events. Detailed operational rules, helper functions, and the handling of infeasible combinations (resulting in the void OIE, $\mathrm{oie}_{\mathrm{void}}$) are provided.
  • Section 5: Algebraic Properties of Sequential Operations – Analyzes the algebraic structure of the two operations, proving properties such as closure, general non-commutativity, and permutation equivalence. The concept of an orbit space for each operation is defined, with the addition operation yielding a single-orbit space and the multiplication operation potentially yielding multiple orbits.
  • Section 6: Implementation of OIE – Describes how an $\mathrm{OIE}$ instance guides the execution of its mapped event. Two types of implementation are defined: one using only the overall interval set ($\mathcal{I}$) and another using the detailed schedule ($\mathcal{F}$) from a sequential operation’s result. A projection mechanism is established to link sequential operations with general $n$-ary finitary operations, explaining how competitions (like 100m dash and downhill skiing) deterministically produce a winner via the permutation-invariant $\min$ operation.
  • Section 7: Full CSA Diagram – Presents the Full Complete Sequential Addition Diagram, a Cayley-table-like representation of the $\oplus$ operation within a finite commutative semigroup featuring an absorbing element. This diagram visually encodes the algebraic properties of CSA and is shown to have correspondences with natural language conjunction (“and”) and programmatic composition.
@@
  • Section 9: Comparison – In the field of computer science, we have conducted comparisons between the content of this paper and process algebra as well as Petri nets in terms of event processing. In the fields of physics and philosophy, we have further drawn comparisons between our interpretation and various physical and philosophical perspectives with respect to the understanding of simultaneous initiation.
@@ 2 Analysis of the properties of event
Event is a noun with complex meanings in both practical and philosophical aspects. It is a fundamental concept frequently used in the fields of physics, computer science, engineering and social sciences (see Casati and Varzi (2023)).
Generally, we refer to the event’s class as “Event”. An instance of Event, as a specific and concrete entity, is represented by the term “event”.
@@ 2.1 Execution status, starting timestamp and ending timestamp of event
Definition 1 (Starting timestamp of an event).
Given an Event instance named event and a timestamp $t_{0}$, if the status of event at the moment $t_{0}$ is running, and for all real numbers $t$ in the set $\{t|t<t_{0}\}$, the status of the event at the corresponding moment is unstarted, then $t_{0}$ is defined as the starting timestamp of event. We denote it as $\mathrm{TS}$. For an instance of $\mathrm{TS}$, we use $\mathrm{ts}$” to represent it.
Definition 2 (Ending timestamp of an event).
Given an Event instance named event and a timestamp $t_{0}^{{}^{\prime}}$, if for all real numbers $t$ in the set $\{t|t\geq t_{0}^{{}^{\prime}}\}$, the status of the event event at the corresponding moment is finished, and there exists a real number $\epsilon>0$ such that for all real numbers $t$ in the set $\{t|t\in(t_{0}^{{}^{\prime}}-\epsilon,t_{0}^{{}^{\prime}})\}$, the status of event at the corresponding moment is running, then $t_{0}^{{}^{\prime}}$ is defined as the ending timestamp of event. We denote it as $\mathrm{TE}$. For an instance of $\mathrm{TE}$, we use $\mathrm{te}$” to represent it.
@@
The naming methods for the noun definitions and instances in Definition 1 and Definition 2 will be used throughout this paper. A noun definition represents what a class of instances is. “What it is” is an abstract level question, while an instance is a specific object created using the noun definition. Therefore, instances are usually used in specific practical actions. For example, consider the following expression:
@@ 2.3 Event with undetermined interval
In this subsection, we introduce the first abstraction of real-world events in this paper, the concept of “Event with Undetermined Interval” (denoted as $\mathrm{Event}^{*}$). This abstraction is achieved by replacing the fixed time interval of an event with undetermined variables, and thereby creating a mapping from the abstract entity back to the original event. While preserving other attributes of the event (such as subject and space) this approach variabilizes the temporal property, enabling flexible assignment or combination of different time intervals in subsequent analysis and operations. Using an example of “A doctor submitting a paper”, we illustrate how a real-world event can be abstracted into an $\mathrm{Event}^{*}$ instance and expressed mathematically, laying the groundwork for the concepts in later chapters.
@@
An abstract entity derived from an Event instance event by specifying the event’s interval in the form of an undetermined variable is defined as an “Event with Undetermined Interval”, denoted as $\mathrm{Event}^{*}$”. And there exists a mapping between $\mathrm{Event}^{*}$ and Event.
$\mathrm{Event}^{*}\rightarrow\textsf{Event}$
Event is typically abstracted to include properties such as subject, time, space, and so on. Definition 5 focuses on variabilizing the property of “interval”; in many cases, we can provide multiple variables for it to choose from.
@@
denoted as $\textsf{event}_{Dr_{submit}}$ the interval of this event is “December 1, 2025, between 10:00 a.m. and 12:00 a.m.”, donated as
@@
We will set this interval as to be determined while keeping all other properties unchanged, and the expression thus becomes follow.
@@
Using definition 5, we can abstract this expression as an $\mathrm{Event}^{*}$ instance $\mathrm{event}^{*}_{Dr_{submit}}$, and
$\mathrm{event}^{*}_{Dr_{submit}}\rightarrow\textsf{event}_{Dr_{submit}}$
$\mathrm{Event}^{*}$ also has atomic and composite types.
@@
An $\mathrm{Event}^{*}$ instance is an atomic $\mathrm{Event}^{*}$ instance if it maps to an AtomEvent instance, denoted as $\textsf{A}\mathrm{Event}^{*}$”.
While an $\mathrm{Event}^{*}$ instance is a composite $\mathrm{Event}^{*}$ instance if it maps to a CompEvent instance, denoted as $\textsf{C}\mathrm{Event}^{*}$”.
@@ 3 Optional intervals event and analysis
Before an Event instance occurs, has someone already arranged for its occurrence? Does the person who arranged it have a thorough understanding of every moment of its execution? If it is possible to arrange an event in advance, then what kind of planning is needed before the event starts to be executed? Is it possible to extract some content from these plans and express it in a mathematical way? This section will revolve around these ideas and propose some new concepts related to events.
Before introducing new concepts in this section, we need to explain some custom characters used in this paper, which are mainly employed to describe several complex mathematical types related to 2-tuples.
$\mathrm{2Tuple}$ $\mathrm{2tuple}$ 2-tuple
@@ 3.1.1 Definition
A class of 4-tuples that map to Event instances and store their available interval information is termed “Optional Intervals Event” (abbreviated as $\mathrm{OIE}$), with the corresponding instance written as $\mathrm{oie}$”.
@@
The 1st element ($\mathcal{C}$) of an $\mathrm{OIE}$ instance is analogous to a bill of components, specifying the subitems that compose the current instance. Its type is a tuple of $\mathrm{OIE}$ instances, which can be an empty tuple “()()”. Some formal examples of $\mathcal{C}$ are provided as follows:
@@
The 2nd element ($\mathcal{F}$) of an $\mathrm{OIE}$ instance is analogous to a detailed schedule-a set whose cardinality corresponds to the number of planned schemes. It records the feasible execution interval combinations of each component for every scheme. It is a $\mathrm{2Tuple}\mathcal{TS}$ instance (a set of tuples of 2-tuples), which can also be $\emptyset$. The form is as follows:
@@
$(\mathrm{ts}_{11},\mathrm{te}_{11})$ is an available interval of the 1st component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan 1, and $(\mathrm{ts}_{21},\mathrm{te}_{21})$ is an available interval of the 2nd component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan 2. Continuing this traversal to the end, $(\mathrm{ts}_{\lambda{n}},\mathrm{te}_{\lambda{n}})$ is an available interval of the $n$th component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan $\lambda$. The number of such plans equals the cardinality of $\mathcal{F}_{\mathrm{oie}}$, i.e., $\left|{\mathcal{F}_{\mathrm{oie}}}\right|$.
The 3rd item ($\mathcal{I}$) of an $\mathrm{OIE}$ instance is analogous to an overall schedule; it is formed by extracting the final start and end timestamps of each scheme to summarize its execution interval, and then integrating the execution intervals of all schemes into one set. It is a $\mathrm{2Tuple}\mathcal{S}$ instance(a set of 2-tuples), which can also be $\emptyset$. When $\mathcal{I}_{\mathrm{oie}}\neq\emptyset$, its form is as follows:
@@
1759740600, 1759741200 and 1759741800 are all timestamps; they are respectively 08:50, 09:00 and 09:10 on November 6, 2025 (Beijing Time). For $\mathrm{oie}_{1}$, suppose there is an Event instance $\textsf{event}_{1}$ in the real world that satisfies
@@
Then, $\textsf{event}_{2}$ has only one feasible interval, which is
@@
$\mathrm{oie}_{\Omega}$ is a special case, whose 3rd element $\mathcal{I}_{\mathrm{oie}_{\Omega}}$ is $\emptyset$. But if there were an Event instance $\textsf{event}_{\Omega}$ mapped to by $\mathrm{oie}_{\Omega}$, we would have no available interval for it to execute.
The 4th element ($\mathcal{A}$) of an $\mathrm{OIE}$ instance is analogous to atomic practical tasks, which are mapped by the atomic components of $\mathcal{C}_{\mathrm{oie}}$. When analyzing the feasibility of a complex event, we typically decompose it—based on real world conditions—into indivisible atomic events. If even one of the atomic events cannot be executed, the entire complex event becomes infeasible. This element serves two main purposes: first, it reinforces the connection between the abstract $\mathrm{OIE}$ instance and the real world Event instance; second, it plays a critical role in the feasibility evaluations involved in the operations discussed in later sections. Some formal examples of the 4th item are as follows:

### v5

@@ Abstract
This paper proposes an algebraic framework for analyzing event execution intervals and sequences, introducing the concept of Optional Intervals Event (OIE) as a core abstraction. The $\mathrm{OIE}$ 4-tuple realizes the mapping of real-world events, storing all feasible execution intervals and sub-event relationships. Based on this, two sequential operations are defined: Complete Sequential Addition ($\oplus$) models concurrent events with a certain degree of equal opportunity within a shared time domain, Complete Sequential Multiplication ($\otimes$) models sequential events with a strict order. And the algebraic properties and orbit spaces of these operations are analyzed.
Addressing the traditional challenge of “simultaneous start”, this paper argues that pointwise temporal equality of timestamps is an unattainable ideal due to physical limits of observational precision. Instead, using sequential operation, “simultaneous start” is reformulated as a certain degree of equal opportunity within an equipotential time window, rather than observed instantaneous equality. This paradigm shift emphasizes integrating subjective algebraic structures to move the perspective from passive observation to active constraint-aware planning, providing a more rigorous formal foundation for event scheduling, physical simulation, and computational modeling that can handle inherent observational uncertainties.
@@ 1. How can we accurately describe the so-called “simultaneous start of two events”?
As of the end of 2024, the error of the most advanced atomic clock in the world is $8\times 10^{-19}$ seconds(see Aeppli et al. (2024)), that is to say, it cannot be guaranteed that the so-called simultaneously starting moments of the two observed events can be accurate to $10^{-1000}$ seconds. So, strictly speaking, this scenario is not truly “at the same time”, we should pursue a more rigorous expression. Can we use an operation *, represented by the following n-ary finitary operation
@@
In a 100 metres race, there are usually 8 athletes competing, it is a head-to-head type of race. In this paper, we use an n-ary finitary operation $\oplus$” to represent a race in which all athletes participate together. Suppose the 8 athletes are arranged according to the lane order, with the $i$-th athlete $athlete_{i}$ in the $i$-th lane. Then a 100 metres race can be expressed as
@@
Swapping the lanes does not affect the starting method of each athlete, they all start running after hearing the starting gun. So intuitively, this operation exhibits a certain kind of transformation invariance.
Downhill skiing is a winter sport. Each skier competes on a track in a certain order one by one, and there can only be one person on the track at a time, so this is a sequential type of race. We use an n-ary finitary operator $\otimes$” to represent a race in which all skiers participate. Then a race with 6 skiers can be expressed as
@@
The skier represented on the left side of parameters starts the race earlier than the skier represented on the right side. Intuitively, this operation does not appear to exhibit transformation invariance.
@@
This paper introduces a paradigm shift from passive observation to constraint-aware subjective planning of events. By leveraging mathematical frameworks of set theory, group theory, and algebraic structures (e.g., Cartesian product, orbit space, permutational equivalence), it establishes a novel formal algebraic methodology for the analysis of event execution, sequence relationships, and object motion.
The subsequent sections are structured as follows:
@@
  • Section 3: Optional Intervals Event (OIE) – Proposes the core abstraction of Optional Intervals Event, 4-tuples $(\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A})$ that map to real-world events. These 4-tuples store all feasible execution intervals and the relationships among sub-events, serving as formal tools for pre-execution planning. The concepts of atomic ($\textsf{Atom}\mathrm{OIE}$) and composite ($\textsf{Comp}\mathrm{OIE}$) $\mathrm{OIE}$ are defined.
  • Section 4: Two Sequential Operations – Introduces two n-ary finitary operations on $\mathrm{OIE}$ instances. Complete Sequential Addition $(\oplus)$ models concurrent events where all participants have a certain degree of equal opportunity to start and end within a specified domain. Complete Sequential Multiplication $(\otimes)$ models strictly ordered sequential events. Detailed operational rules, helper functions, and the handling of infeasible combinations (resulting in the void $\mathrm{OIE}$ instance $\mathrm{oie}_{\mathrm{void}}$) are provided.
  • Section 5: Algebraic Properties of Sequential Operations – Analyzes the algebraic structure of the two operations, proving properties such as closure, general non-commutativity, and permutation equivalence. The orbit space for each operation is defined, for a finite set of $\mathrm{OIE}$ instances complete sequential addition($\oplus$) yielding a single-orbit space due to permutational equivalence while complete sequential multiplication($\otimes$) - constrained by strict sequential order - potentially yielding multiple orbits.
  • Section 6: Implementation of OIE – Describes how $\mathrm{OIE}$ instances guide the execution of their mapped events. Two types of implementation are defined: one using only the overall interval set ($\mathcal{I}$) and another using the detailed schedule ($\mathcal{F}$) from a sequential operation’s result. A projection mechanism is established to link sequential operations with general $n$-ary finitary operations, explaining how competitions (like 100m dash and downhill skiing) deterministically produce a winner via the permutation-invariant $\min$ operation. In a computer science application, we map the stage structure of Merge Sort to a combination of algebraic operations with distinct commutativity properties for the first time, and analyze the feasibility of parallel sorting algebraically. In an example of probability theory, we have analyzed the use of $\mathrm{OIE}$ in classical probability, providing a more rigorous description for drawing lots.
  • Section 7: Full CSA Diagram – Presents the Full Complete Sequential Addition Diagram, a Cayley-table-like representation of the $\oplus$ operation within a finite commutative semigroup featuring an absorbing element. This diagram visually encodes the algebraic properties of complete sequential addition and is shown to have correspondences with natural language conjunction (“and”) and programmatic composition.
@@
  • Section 9: Comparison – In the field of computer science, we have conducted comparisons between the content of this paper and classical algebraic and formal models for event processing. In the fields of physics and philosophy, we have further drawn comparisons between our interpretation and various physical and philosophical perspectives with respect to the understanding of simultaneous initiation.
@@
  • Appendix: Simulation Code and Diagram Generation Code We provide the urls of simulation code and operation table code.
@@ 2 Analysis of the properties of event
Event is a noun with complex meanings in both practical and philosophical aspects. It is a fundamental concept frequently used in the fields of physics, computer science, engineering and social sciences (see Casati and Varzi (2023)). Generally, we refer to the event’s class as “Event”. An instance of Event, as a specific and concrete entity, is represented by the term “event”. In this Section, we discuss some basic properties of Event.
Subsection 2.1 clarifies the three execution states of an event, the definitions and core properties of starting and ending timestamps, and standardizes the representation form of time intervals.
Subsection 2.2 defines the concepts of atomic events and composite events, and provides the calculation method for the start and end timestamps of composite events.
Subsection 2.3 proposes the abstract concept of Event with Undetermined Interval ($\mathrm{Event}^{*}$), realizes flexible mapping of real-world events by replacing fixed time intervals with undetermined variables, and distinguishes its atomic and composite types.
@@ 2.1 Execution status, starting timestamp and ending timestamp of event
Definition 1 (Starting and ending timestamp of an event).
Given an Event instance named event.
If the status of event at the timestamp $t_{0}$ is running, and for all real numbers $t$ in the set $\{t|t<t_{0}\}$, the status of the event at the corresponding moment is unstarted, then $t_{0}$ is defined as the “starting timestamp” of event. We denote it as $\mathrm{TS}$. For an instance of $\mathrm{TS}$, we use $\mathrm{ts}$” to represent it.
If for all real numbers $t$ in the set $\{t|t\geq t_{0}^{{}^{\prime}}\}$, the status of the event event at the corresponding moment is finished, and there exists a real number $\epsilon>0$ such that for all real numbers $t$ in the set $\{t|t\in(t_{0}^{{}^{\prime}}-\epsilon,t_{0}^{{}^{\prime}})\}$, the status of event at the corresponding moment is running, then $t_{0}^{{}^{\prime}}$ is defined as the “ending timestamp” of event. We denote it as $\mathrm{TE}$. For an instance of $\mathrm{TE}$, we use $\mathrm{te}$” to represent it.
@@
The naming methods for the noun definitions and instances in Definition 1 will be used throughout this paper. A noun definition represents what a class of instances is. “What it is” is an abstract level question, while an instance is a specific object created using the noun definition. Therefore, instances are usually used in specific practical actions. For example, consider the following expression:
@@ 2.3 Event with undetermined interval
In this subsection, we introduce the first abstraction of real-world events in this paper, the concept of “Event with Undetermined Interval” (denoted as $\mathrm{Event}^{*}$). This abstraction is achieved by replacing the fixed time interval of an event with undetermined variables, and thereby creating a bijective mapping between the abstract entity back and the original event. While preserving other attributes of the event (such as subject and space) this approach variabilizes the temporal property, enabling flexible assignment or combination of different time intervals in subsequent analysis and operations. Using an example of “A doctor submitting a paper”, we illustrate how a real-world event can be abstracted into an $\mathrm{Event}^{*}$ instance and expressed mathematically, laying the groundwork for the concepts in later chapters.
@@
An abstract entity derived from an Event instance event by specifying the event’s interval in the form of an undetermined variable is defined as an “Event with Undetermined Interval”, denoted as $\mathrm{Event}^{*}$”. And there exists a bijective mapping between $\mathrm{Event}^{*}$ and Event, denoted as
$\mathrm{Event}^{*}\leftrightarrow\textsf{Event}$
Event is typically abstracted to include properties such as subject, space, etc. In particular, it should be emphasized that an event has a clear and fixed interval Definition 4 precisely focuses on variabilizing the property of “interval”. In many cases, we can provide multiple variables for it to choose from.
@@
which denoted as $\textsf{event}_{Dr}$, the interval of this event is “December 1, 2025, between 10:00 a.m. and 12:00 a.m.”, donated as
@@
If we set this interval as “to be determined” while keeping all other properties unchanged, the expression thus becomes as following
@@
Using definition 4, we can abstract this expression as an $\mathrm{Event}^{*}$ instance $\mathrm{event}^{*}_{Dr}$, and
$\mathrm{event}^{*}_{Dr}\leftrightarrow\textsf{event}_{Dr}$
@@
An $\mathrm{Event}^{*}$ instance is an atomic $\mathrm{Event}^{*}$ instance if it bijectively maps to an AtomEvent instance, denoted as $\textsf{A}\mathrm{Event}^{*}$”; while an $\mathrm{Event}^{*}$ instance is a composite $\mathrm{Event}^{*}$ instance if it bijectively maps to a CompEvent instance, denoted as $\textsf{C}\mathrm{Event}^{*}$”.
@@ 3 Optional intervals event and analysis
Before an event occurs, is it prearranged? Does the planner fully grasp every moment of its execution? What pre-execution planning is required for such arranged events, and can we extract and mathematically formalize key elements from these plans? This section explores these questions and introduces novel event-related concepts.
Section 3.1 formally defines Optional Intervals Event ($\mathrm{OIE}$) as a 4-tuple ($\mathcal{C},\mathcal{F},\mathcal{I},\mathcal{A}$) that maps to real-world events and stores feasible execution intervals and sub-event relationships, along with its properties and classifications ($\textsf{Atom}\mathrm{OIE}$$\textsf{Comp}\mathrm{OIE}$).
Section 3.2 introduces $\mathrm{OIE}\mathcal{S}$(a set of OIE instances) and focuses on the set of feasible interval 2-tuple sets derived from its members.
Section 3.3 defines the void OIE instance ($\mathrm{OIE}_{\mathrm{void}}$) which cannot map to any real-world event and serves as the result of infeasible sequential operations.
To clarify the description of complex mathematical types involving 2-tuples, the following conventions are adopted (“Type Name”: abstract category with capitalized first letter; “Type Instance”: specific object with lowercase first letter):
Type Name Type Instance Core Meaning
$\mathrm{2Tuple}$ $\mathrm{2tuple}$ 2-tuple (basic core concept, a single 2-tuple)
@@ 3.1.1 Definition
A class of 4-tuples that map to Event instances and store their available interval information, termed “Optional Intervals Event” (abbreviated as $\mathrm{OIE}$), with the corresponding instance written as $\mathrm{oie}$”.
@@
The 1st element ($\mathcal{C}$) of an $\mathrm{OIE}$ instance is analogous to “a bill of components”, specifying the subitems that compose the current instance. Its type is a tuple of $\mathrm{OIE}$ instances, which can be an empty tuple “()()”. Some formal examples of $\mathcal{C}$ are provided as follows:
@@
The 2nd element ($\mathcal{F}$) of an $\mathrm{OIE}$ instance is analogous to “a detailed schedule set” whose cardinality corresponds to the number of planned schemes. It records the feasible execution interval combinations of each component for every scheme. It is a $\mathrm{2Tuple}\mathcal{TS}$ instance (a set of tuples of 2-tuples), which can also be $\emptyset$. Its form is as follows:
@@
$(\mathrm{ts}_{11},\mathrm{te}_{11})$ is an available interval of the 1st component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan 1, and $(\mathrm{ts}_{21},\mathrm{te}_{21})$ is an available interval of the 1st component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan 2. Continuing this traversal to the end, $(\mathrm{ts}_{\lambda{n}},\mathrm{te}_{\lambda{n}})$ is an available interval of the $n$th component $\mathrm{OIE}$ instance in $\mathcal{C}_{\mathrm{oie}}$ under schedule plan $\lambda$, and it satisfies $\mathrm{ts}_{\lambda{n}}<\mathrm{te}_{\lambda{n}}$. The number of such plans equals the cardinality of $\mathcal{F}_{\mathrm{oie}}$, i.e., $\left|{\mathcal{F}_{\mathrm{oie}}}\right|$.
The 3rd item ($\mathcal{I}$) of an $\mathrm{OIE}$ instance is analogous to “an overall schedule”, it is formed by extracting the final start and end timestamps of each scheme to summarize its execution interval, and then integrating the execution intervals of all schemes into one set. It is a $\mathrm{2Tuple}\mathcal{S}$ instance(a set of 2-tuples), which can also be $\emptyset$. When $\mathcal{I}_{\mathrm{oie}}\neq\emptyset$, its form is as follows:
@@
1759740600, 1759741200 and 1759741800 are all timestamps, they are respectively 08:50, 09:00 and 09:10 on November 6, 2025 (Beijing Time). For $\mathrm{oie}_{1}$, suppose there is an Event instance $\textsf{event}_{1}$ in the real world that satisfies
@@
and $\textsf{event}_{2}$ has only one feasible interval, which is
@@
$\mathrm{oie}_{\Omega}$ is a special case, whose 3rd element $\mathcal{I}_{\mathrm{oie}_{\Omega}}$ is $\emptyset$. However, if there were an Event instance $\textsf{event}_{\Omega}$ mapped to by $\mathrm{oie}_{\Omega}$, we would have no available interval for it to execute.
The 4th element ($\mathcal{A}$) of an $\mathrm{OIE}$ instance is analogous to “atomic practical tasks”, which are mapped by the atomic components of elements in $\mathcal{C}_{\mathrm{oie}}$. When analyzing the feasibility of a complex event, we typically decompose it—based on real world conditions—into indivisible atomic events. If even one of the atomic events cannot be executed, the entire complex event becomes infeasible. This element serves two main purposes: first, it reinforces the connection between the abstract $\mathrm{OIE}$ instance and the real world Event instance; second, it plays a critical role in the feasibility evaluations involved in the operations discussed in later sections. Some formal examples of the 4th item are as follows: